OpenStudy (anonymous):

An angle that has a supplement also has a complement. Always, sometimes, or never?

hartnn (hartnn):

Let that angle be x. it has supplement 180-x and complement 90-x. so this can ALWAYS be true.

OpenStudy (anonymous):

But wouldn't it have a limit of 180 degrees? Or does it just go all around?

hartnn (hartnn):

it just go all around. let x=200 supplement = 180-200 = -20 = 340 complement =90-200=-110=360-110=250.

OpenStudy (anonymous):

I'm still lost are there any pictures you can draw because all the equations just confused me even more.

hartnn (hartnn):